Find an equation of the tangent line to the curve 9x³ – y³ = 1 at the point (0, -1).

1
Expert's answer
2021-12-06T17:10:59-0500

Let us find an equation of the tangent line to the curve "9x^3 \u2013 y^3 = 1" at the point "(0, -1)." It follows that "27x^2 \u2013 3y^2y' = 0," and hence "y'(0)=\frac{27\cdot 0^2}{3(-1)^2}=0." Using the formula "y=y_0+y'(x_0)(x-x_0)" for "x_0=0" and "y_0=-1," we conclude that "y=-1+0(x-0)=-1", that is "y=-1" is an equation of the tangent line to the curve "9x^3 \u2013 y^3 = 1" at the point "(0, -1)."